OpenStudy (anonymous):

Find the value of: 1. 10√2 over √5 2. √8 over 9 3. 3x over √x^3 4. √3 over 3 5. 2√3 over √10 6. √6 over 5√3

OpenStudy (radar):

\[10\sqrt{2}\over \sqrt{5}\] should be "rationalized", that is eliminate the radical in the denominator. Are you familiar with that procedure?

OpenStudy (anonymous):

No. :/

OpenStudy (radar):

You multiply both the numerator and denominator by the radical in the denominator. In this case it will look like this:\[10\sqrt{2}\sqrt{5}\over \sqrt{5}\sqrt{5}\]

OpenStudy (radar):

\[10\sqrt{2}\sqrt{5}\over5\]\[2\sqrt{10}\]Pretty simple huh.

OpenStudy (anonymous):

A lil. Thanks a bunch though. :)

OpenStudy (radar):

Lets do 6.\[\sqrt{6}\over5\sqrt{3}\]

OpenStudy (anonymous):

Okay. :)

OpenStudy (anonymous):

I think I get it... is number 2: 2√2 over 3?

OpenStudy (radar):

\[1/5 \times \sqrt{6}/\sqrt{3}=1/5 \times \sqrt{2}\]or\[\sqrt{2}\over5\]or\[0.2\sqrt{2}\]

OpenStudy (radar):

2.\[2\sqrt{2}\over9 \]provided the radical did not include the 9. If the problem is\[\sqrt{8/9}\] then it is \[2\sqrt{2}\over3\] I am not sure which one you mean for #2.

OpenStudy (radar):

Good luck with these things, I have to go back to work.(yard work)lol

OpenStudy (anonymous):

thank you so much. :)

OpenStudy (radar):

Which one did you mean for #2, you worked it out as if 9 was included in the radical was that way it was meant?

OpenStudy (anonymous):

yes it was.